OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6? A. Y = 2x+6 B. Y = -2/3x + 5/3 C. Y = 2x + 3 D. Y = 2/3x + 5/3

OpenStudy (luigi0210):

Put the equation in slope intercept form.

OpenStudy (anonymous):

Y=mx+b?

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

y=2x+3 i think?

OpenStudy (luigi0210):

Not exactly

OpenStudy (anonymous):

Im on a timed test can I get the answer please? Im about to fail if I dont get this

OpenStudy (luigi0210):

Okay, so we have this: \[\large 2x+3y=6\] And we need to get it to look like this: \[\Large y=mx+b\] So we need to get y by itself. We subtract 2x from both sides: \[\large 2x-2x+3y=6-2x\] Simpliffy and we get this: \[\large 3y=6-2x\] Now we just divide everything by 3, and we end up with: \[\large y=2-\frac{2}{3}x\] Or in slope intercept form: \[\large y=-\frac{2}{3}x+2\]

OpenStudy (luigi0210):

make sense?